For calibration of 10 ml at 23 degree centigrade a 10 ml liquor of water is weighted by analytical balance which is 9.9367 g. Calculate the actual transferred volume. If the density of water at 23 degree centigrade is 0.99754 g/ ml.
1
Expert's answer
2016-12-13T11:19:09-0500
ρ=m/V

V=m/ρ

V=9.9367/0.99754=9.9612 ml